OpenStudy (anonymous):

What is 32/36 in simplest form? 16/18 4/9 5/6 8/9

OpenStudy (anonymous):

Divide both numbers by 4

OpenStudy (legends):

Yea @TheDream

OpenStudy (phi):

btw, there are "tricks" to find what number divides into another. see https://en.wikipedia.org/wiki/Divisibility_rule#Divisibility_rules_for_numbers_1.E2.80.9320 I would learn at least the first few. in your problem we could do it in steps. both top and bottom are even, so both are divisible by 2. we then have 32/36 = 16/18 but 16 and 18 are both divisible by 2, so simplify again: 16/18= 8/9 8 is 2*2*2 9 is 3*3 they cannot be simplified any more (we can't divide the same number into both 8 and 9( so 8/9 is the answer